# Zeal: Rethinking Large-Scale Resource Allocation with “Decouple and Decompose”

### v1

@@ Abstract
Resource allocation is fundamental for cloud systems to ensure efficient resource sharing among tenants. However, the scale of such optimization problems has outgrown the capabilities of commercial solvers traditionally employed in production. To scale up resource allocation, prior approaches either tailor solutions to specific problems or rely on assumptions tied to particular workloads. In this work, we revisit real-world resource allocation problems and uncover a common underlying structure: a vast majority of these problems are inherently separable, i.e., they optimize the aggregate utility of individual resource and demand allocations, under separate constraints for each resource and each demand. Building on this insight, we develop Zeal, a general, scalable, and theoretically grounded framework for accelerating resource allocation through a “decouple and decompose” approach. Zeal systematically decouples entangled resource and demand constraints, thereby decomposing the overall optimization into alternating per-resource and per-demand allocations, which can then be solved efficiently and in parallel. We have implemented Zeal as a library extension to an open-source solver, maintaining a familiar user interface. Experimental results across three prominent resource allocation tasks—traffic engineering, cluster scheduling, and load balancing—demonstrate Zeal’s substantial speedups and robust allocation quality.
@@ 1 Introduction
In this study, we re-examine real-world resource allocation problems through a different lens and aim to develop a general and scalable optimization framework without catering to specific problems or making workload-dependent assumptions. Our approach, Zeal, is theoretically grounded and enables fine-grained decomposition of the optimization problem, alternating between per-resource and per-demand allocations that can be solved in parallel (Figure 1).
Figure 1: Overview of Zeal.
@@
Given the intertwined nature of resource and demand constraints, Zeal first decouples them as follows. Let $x$ denote the resource allocation matrix between $n$ resources and $m$ demands, where $x_{ij}$ represents the amount of resource $i$ assigned to demand $j$. The key mathematical trick in Zeal is to introduce a duplicate of the matrix $x$, denoted as $z$, along with an additional constraint $x=z$. Next, Zeal iteratively and alternately solves for $x$ subject only to resource constraints with $z$ fixed, and solves for $z$ subject only to demand constraints with $x$ fixed. In each iteration, Zeal further ensures that $x$ and $z$ do not deviate significantly from each other in order to satisfy $x=z$. This decoupling transformation lends itself to a classical constrained optimization framework called ADMM (Alternating Direction Method of Multipliers) [9]. ADMM allows Zeal to embed each of the resource and demand constraints into the original objective as a penalty for constraint violation, thereby constructing augmented Lagrangians [23] as new optimization objectives. ADMM provides the theoretical foundation that enables alternating optimizations of these augmented Lagrangians with respect to $x$ and $z$, respectively.
@@
Next, Zeal further decomposes the above optimizations into per-resource and per-demand allocations. This is feasible owing to the separable structure we identified across resource allocation problems: the original objective is separable on each resource and demand variable, and so are the penalty terms converted by ADMM from constraints in the augmented Lagrangians. After decomposition, per-resource allocations involve solving $n$ subproblems, each with $m$ demand variables, and $m$ subproblems with $n$ resource variables for per-demand allocations. Consequently, the scale of optimization is greatly reduced from $n\times m$ variables to only $n$ or $m$ variables in each subproblem. These subproblems can be solved independently with massive parallelism using off-the-shelf solvers. Figure 2 illustrates the “decouple and decompose” workflow of Zeal.
@@
To demonstrate Zeal’s practicality, we have implemented Zeal as a Python package [^3] Zeal can be installed with pip install zeal after publication., built on the widely used open-source solver cvxpy [17, 4]. The interface of Zeal is also designed to resemble cvxpy for ease of use (§6). Moreover, instead of simulating the parallel execution of subproblems on multiple CPU cores as previous studies did [1, 43], we have carefully optimized Zeal’s implementation using Ray [40], enabling efficient utilization of multiple cores without incurring significant overheads, such as cache contention.
@@
We evaluate Zeal in three prominent resource allocation problems—traffic engineering, cluster scheduling, and load balancing—and demonstrate its faster speed and higher allocation quality than the state of the art (§7). Compared with the best variant of POP [43] for each problem, Zeal achieves a 5.3% higher allocation quality and a 7.6$\times$ speedup in traffic engineering, 14% and 1.34$\times$ in cluster scheduling, and 5.3% and 1.13$\times$ in load balancing.
Figure 2: Workflow of Zeal.
@@ 2 Real-World Resource Allocation Problems
3 Zeal: Decouple and Decompose
We present Zeal, a general, scalable, and theoretically rooted framework for optimizing real-world resource allocation problems. Leveraging the separable structure (§2) that underpins a vast majority of these problems, Zeal enables massive parallelism with near-optimal solutions through a technique that we call “decouple and decompose.” A formal, mathematical representation of Zeal is included in Appendix A.
@@ 3.2 Decompose
thereby decomposing the allocation of $x$ with $n\times m$ variables into $n$ per-resource allocations of $x_{i*}$, each with $m$ variables. To solve them, Zeal uses an off-the-shelf solver (cvxpy in our implementation) to find (exact or approximate) solutions while making liberal use of all available CPU cores.
@@ 4 Generality and Limitations
In this section, we start with the generality of Zeal, and then discuss its limitations, including scenarios with limited parallelism and chances of suboptimal solutions.
@@ 4.1 Generality
Zeal can be applied to a wide range of resource allocation problems. Table 1 presents a summary of real-world resource allocation problems in the recent literature, mostly from OSDI, SOSP, NSDI, and SIGCOMM in the past four years. Next, we explain why Zeal is applicable.
  • Variables.  Inheriting from ADMM, Zeal inherently supports real-number (float) allocation variables. Empirical evidence[53] and theoretical proofs [56, 38, 60] on ADMM also indicate that Zeal can effectively handle boolean and integer variables by projecting real-number solutions to desired data types over iterations.
  • Objective.  Zeal requires a max or sum of convex utility functions over per-resource or per-demand allocations. Most objectives used in systems exhibit convexity, such as the weighted allocation sum [44, 63], the logarithmic resource utility [3], and the quadratic allocation cost [64, 59].
  • Constraints.  Zeal assumes the linearity of resource and demand constraints, which holds in all our surveyed problems, e.g., the sum of demands allocated to a resource must not exceed its capacity. The constraints in these problems are per resource or per demand as well. Even under a constraint on multiple resources or demands, Zeal can be relaxed to groups of resources or demands. Zeal bundles these groups together into per-resource-group or per-demand-group allocations during decomposition.
Unlike POP [43], Zeal works with any workload, not only “granular” ones. In POP, the demands in each decomposed problem can access only a subset of resources, assuming each demand only requests a small fraction of interchangeable resources such that a resource subset suffices. This granular assumption can be unrealistic in practice, due to uneven demand distributions [28, 57] or limited resources that each demand is allowed to utilize [58, 33]. In contrast, each decomposed problem in Zeal can still access all the resources or demands; thus, the granular assumption is no longer required, affirming Zeal’s generality to various workload types.
@@ 4.2 Limitations
While Zeal applies to a variety of resource allocation problems, it might not always be the most suitable solution for every allocation problem.
Limited parallelism.  Zeal achieves parallelism by decomposing the full problem into per-resource and per-demand allocations. Therefore, when we cannot fully decompose the problem, Zeal will experience limited parallelism, although this does not happen in our survey (Table 1).
@@
Suboptimal solutions.  It is possible for Zeal not to attain an ideal solution when the allocation problem deviates from the standard ADMM formulation.
  • Non-convex problems.  ADMM can solve non-convex optimization problems [53, 62, 68] empirically, providing theoretical guarantees only under specific assumptions[56, 38, 60]. As a result, Zeal can output suboptimal solutions in certain non-convex cases.
@@ 5 Case Studies
In this section, we will present case studies on how Zeal can be applied to three resource allocation problems: traffic engineering, cluster scheduling, and load balancing, as also studied in POP [43].
@@ 5.1 Traffic Engineering
  • Objective.  The link utilization of a link $(u,v)$ is the ratio between the total flow allocated on that link and link capacity. Our objective is to minimize the maximum utilization of each link:
@@
Using Zeal.  The traffic engineering problem aligns the separable structure for Zeal: the objectives are the sum of allocation quality per demand or the maximum of allocation quality per link, respectively. Linear constraints are applied to each demand and each link. Therefore, Zeal can decompose the problem into $|E|$ per-link allocations and $|V|^{2}$ per-demand allocations for parallelism.
@@ 5.2 Cluster Scheduling
Using Zeal.  The cluster scheduling exhibits the separable structure required by Zeal: The objective function is a minimum of allocation quality per job and a sum of allocation quality per job, respectively. Additionally, linear constraints are applied to each job and each computing resource. Thus, Zeal can decompose the problem into $n$ per-job allocations and $m$ per-demand allocations.
@@ 5.3 Load Balancing
Using Zeal.  Given the separable structure, where the objective is the sum of allocation quality per data shard and linear constraints apply to each server, and each shard, Zeal can decompose the problem into $n$ per-server allocations and $m$ per-shard allocations.
6 Implementation of Zeal
We implement Zeal as a Python package. After publication, we will upload Zeal to PyPI for easy installation (pip install zeal).
Listing 1 showcases a simple resource allocation problem solved using our Zeal package. In line 4, we initially create a $N\times M$ non-negative allocation variable x, where x[i,j] represents the allocation of resource i to demand j. We then create a constraint for each resource in line 7–8 and a constraint for each demand in line 9–10, as well as an objective of minimizing the sum of x in line 13. Finally, we construct the problem in line 16–17 and solve it in line 18.
1import zeal
@@ 6 Implementation of Zeal
3# Create allocation variables.
4x = zeal.Variable((N, M), nonneg=True)
@@
12# Create an objective.
13obj = zeal.Minimize(x.sum())
@@
15# Construct and solve the problem.
16prob = zeal.Problem(
@@
Listing 1: Resource allocation example with Zeal package.
Zeal borrows the interface from cvxpy [17, 4]. Zeal inherits most of cvxpy’s methods, e.g. Variable(.) in line 4, Minimize(.) in line 13. One difference from cvxpy is that users must specify resource constraints and demands constraints separately (line 16–17) when initializing a problem. As a parallel solution, Zeal allows users to specify the number of CPU cores (line 18); if not specified, Zeal uses all available CPU cores.
Behind Zeal’s interface, the resource allocation problem is solved in the following three steps.
@@ 7 Evaluation
    How does Zeal compare with state-of-the-art approaches in allocation quality and computation time? (§7.1)
@@
    How do Zeal and state-of-the-art approaches react to changes in problem granularity, temporal fluctuation, spatial redistribution, and link failures? (§7.2)
@@
    What is the contribution of each design in Zeal? (§7.3)
@@
  • Zeal.  Zeal is our truly parallel implementation, with details described in Section 6.
  • Zeal.  To facilitate a fair comparison with POP, we create a variant of Zeal, denoted as Zeal, following POP-style simulation where subproblems are solved sequentially and a parallel solving time is calculated mathematically.
@@
7.1 Zeal vs. the State of the Art
In this section, we compare Zeal against state-of-the-art approaches on resource allocation problems from Section 5.
@@ 7.1.1 Traffic Engineering
Maximize total flow.  Figure 3 compares Zeal to state-of-the-art methods in maximizing total flow. Zeal achieves 90.8% satisfied demand within 30 seconds and 92% within 60 seconds, demonstrating a superior trade-off between satisfied demand and computation time compared to Exact Sol, POP, and Pinning. While POP-4 also reaches 92% satisfied demand, its average computation time is significantly longer at 1658 seconds. POP-16 and POP-64 improve parallelism by dividing the workload into more subproblems, reducing computation times to 456 seconds and 380 seconds, respectively. Note that the speedup of POP-64 is limited compared to POP-16 as the number of cores per subproblem drops from 4 to 1. This increased parallelism also comes at a cost, with satisfied demand decreasing to 87.4% for POP-16 and 81.6% for POP-64. Pinning focuses on top demands to reduce problem size but remains constrained by the sequential nature of its optimization solvers. As a result, it achieves a modest speedup, requiring 149 seconds to reach 87.3% satisfied demand. Teal, a learning-accelerated approach, leverages massive GPU parallelism to achieve 89% satisfied demand in just 1 second. However, it relies on a highly customized machine-learning framework, demanding significant human effort for design and training.
Zeal, representing the same algorithm under ideal conditions as POP, achieves a satisfied demand that is 8% higher than Zeal within 10 seconds in Figure 3 and consistently delivers superior allocation quality across the other figures. The reason is that Zeal executes each iteration more quickly, allowing it to complete more iterations within a given time, resulting in higher allocation quality. The difference in iteration time between Zeal and Zeal arises from three factors, as discussed in more detail in § 7.3: (1) Zeal measures the end-to-end time, including problem-solving, compilation, solution unpacking, and other associated tasks, while Zeal only takes the problem-solving time into account. (2) Zeal’s parallel implementation (§ 6) incurs overhead from cache contention, which slows down the problem-solving process. (3) In Zeal, each subproblem is statically pre-assigned to one of the processes, making it susceptible to stragglers. In contrast, Zeal assumes perfect dynamic scheduling.
@@
Min-max link utilization.  Figure 4 compares Zeal with other approaches in minimizing maximum link utilization. Zeal achieves a low maximum link utilization of 1.67 within 10 seconds and 1.71 within 5 seconds. Exact Sol reaches 1.63 with an average computation time of 35 seconds, and it is faster than itself in maximizing total flow because minimizing max link utilization is "easier" and requires fewer iterations. Pinning takes 42.5 seconds, making it slower than Zeal and slightly slower than Exact Sol, as it requires additional time to rebuild the problem when the top 10% of node pairs change. POP’s variants, POP-4, POP-16, and POP-64, achieve a maximum link utilization of 1.70, 1.77, and 1.95 with 9, 2, and 0.4 seconds on average, respectively. Although POP-16 is faster than Zeal, a fairer comparison between POP-16 and Zeal, both simulating parallelism, reveals that Zeal achieves a comparable computation time to POP-16. The domain-specific solution Teal takes only 0.3 seconds with massive parallelism from GPU with a maximum link utilization of 1.74.

### v2

# Decouple and Decompose: Scaling Resource Allocation with DeDe

@@ Abstract
Resource allocation is fundamental for cloud systems to ensure efficient resource sharing among tenants. However, the scale of such optimization problems has outgrown the capabilities of commercial solvers traditionally employed in production. To scale up resource allocation, prior approaches either tailor solutions to specific problems or rely on assumptions tied to particular workloads. In this work, we revisit real-world resource allocation problems and uncover a common underlying structure: a vast majority of these problems are inherently separable, i.e., they optimize the aggregate utility of individual resource and demand allocations, under separate constraints for each resource and each demand. Building on this insight, we develop DeDe, a general, scalable, and theoretically grounded framework for accelerating resource allocation through a decouple-and-decompose approach. DeDe systematically decouples entangled resource and demand constraints, thereby decomposing the overall optimization into alternating per-resource and per-demand allocations, which can then be solved efficiently and in parallel. We have implemented DeDe as a library extension to an open-source solver, maintaining a familiar user interface. Experimental results across three prominent resource allocation tasks—traffic engineering, cluster scheduling, and load balancing—demonstrate DeDe’s substantial speedups and robust allocation quality.
@@ 1 Introduction
In this study, we re-examine real-world resource allocation problems through a different lens and aim to develop a general and scalable optimization framework without catering to specific problems or making workload-dependent assumptions. Our approach, DeDe, is theoretically grounded and enables fine-grained decomposition of the optimization problem, alternating between per-resource and per-demand allocations that can be solved in parallel (Figure 1).
Figure 1: Overview of DeDe.
@@
Given the intertwined nature of resource and demand constraints, DeDe first decouples them as follows. Let $x$ denote the resource allocation matrix between $n$ resources and $m$ demands, where $x_{ij}$ represents the amount of resource $i$ assigned to demand $j$. The key mathematical trick in DeDe is to introduce a duplicate of the matrix $x$, denoted as $z$, along with an additional constraint $x=z$. Next, DeDe iteratively and alternately solves for $x$ subject only to resource constraints with $z$ fixed, and solves for $z$ subject only to demand constraints with $x$ fixed. In each iteration, DeDe further ensures that $x$ and $z$ do not deviate significantly from each other in order to satisfy $x=z$. This decoupling transformation lends itself to a classical constrained optimization framework called ADMM (Alternating Direction Method of Multipliers) [9]. ADMM allows DeDe to embed each of the resource and demand constraints into the original objective as a penalty for constraint violation, thereby constructing augmented Lagrangians [23] as new optimization objectives. ADMM provides the theoretical foundation that enables alternating optimizations of these augmented Lagrangians with respect to $x$ and $z$, respectively.
@@
Next, DeDe further decomposes the above optimizations into per-resource and per-demand allocations. This is feasible owing to the separable structure we identified across resource allocation problems: the original objective is separable on each resource and demand variable, and so are the penalty terms converted by ADMM from constraints in the augmented Lagrangians. After decomposition, per-resource allocations involve solving $n$ subproblems, each with $m$ demand variables, and $m$ subproblems with $n$ resource variables for per-demand allocations. Consequently, the scale of optimization is greatly reduced from $n\times m$ variables to only $n$ or $m$ variables in each subproblem. These subproblems can be solved independently with massive parallelism using off-the-shelf solvers. Figure 2 illustrates the “decouple and decompose” workflow of DeDe.
@@
To demonstrate DeDe’s practicality, we have implemented it as a Python package, installable via pip install dede. DeDe is built on the widely used open-source solver cvxpy, and its interface is also designed to resemble cvxpy for ease of use (§6). Moreover, instead of simulating the parallel execution of subproblems on multiple CPU cores as previous studies did [1, 43], we have carefully optimized DeDe’s implementation using Ray [40], enabling efficient utilization of multiple cores without incurring significant overheads, such as cache contention.
@@
We evaluate DeDe in three prominent resource allocation problems—traffic engineering, cluster scheduling, and load balancing—and demonstrate its faster speed and higher allocation quality than the state of the art (§7). Compared with the best variant of POP [43] for each problem, DeDe achieves a 5.3% higher allocation quality and a 7.6$\times$ speedup in traffic engineering, 14% and 1.34$\times$ in cluster scheduling, and 5.3% and 1.13$\times$ in load balancing.
Figure 2: Workflow of DeDe.
@@ 2 Real-World Resource Allocation Problems
3 DeDe: Decouple and Decompose
We present DeDe, a general, scalable, and theoretically rooted framework for optimizing real-world resource allocation problems. Leveraging the separable structure (§2) that underpins a vast majority of these problems, DeDe enables massive parallelism with near-optimal solutions through a decouple-and-decompose technique. A formal, mathematical representation of DeDe is included in Appendix A.
@@ 3.2 Decompose
thereby decomposing the allocation of $x$ with $n\times m$ variables into $n$ per-resource allocations of $x_{i*}$, each with $m$ variables. To solve them, DeDe uses an off-the-shelf solver (cvxpy in our implementation) to find (exact or approximate) solutions while making liberal use of all available CPU cores.
@@ 4 Generality and Limitations
In this section, we start with the generality of DeDe, and then discuss its limitations, including scenarios with limited parallelism and chances of suboptimal solutions.
@@ 4.1 Generality
DeDe can be applied to a wide range of resource allocation problems. Table 1 presents a summary of real-world resource allocation problems in the recent literature, mostly from OSDI, SOSP, NSDI, and SIGCOMM in the past four years. Next, we explain why DeDe is applicable.
  • Variables.  Inheriting from ADMM, DeDe inherently supports real-number (float) allocation variables. Empirical evidence[53] and theoretical proofs [56, 38, 60] on ADMM also indicate that DeDe can effectively handle boolean and integer variables by projecting real-number solutions to desired data types over iterations.
  • Objective.  DeDe requires a max or sum of convex utility functions over per-resource or per-demand allocations. Most objectives used in systems exhibit convexity, such as the weighted allocation sum [44, 63], the logarithmic resource utility [3], and the quadratic allocation cost [64, 59].
  • Constraints.  DeDe assumes the linearity of resource and demand constraints, which holds in all our surveyed problems, e.g., the sum of demands allocated to a resource must not exceed its capacity. The constraints in these problems are per resource or per demand as well. Even under a constraint on multiple resources or demands, DeDe can be relaxed to groups of resources or demands. DeDe bundles these groups together into per-resource-group or per-demand-group allocations during decomposition.
Unlike POP [43], DeDe works with any workload, not only “granular” ones. In POP, the demands in each decomposed problem can access only a subset of resources, assuming each demand only requests a small fraction of interchangeable resources such that a resource subset suffices. This granular assumption can be unrealistic in practice, due to uneven demand distributions [28, 57] or limited resources that each demand is allowed to utilize [58, 33]. In contrast, each decomposed problem in DeDe can still access all the resources or demands; thus, the granular assumption is no longer required, affirming DeDe’s generality to various workload types.
@@ 4.2 Limitations
While DeDe applies to a variety of resource allocation problems, it might not always be the most suitable solution for every allocation problem.
Limited parallelism.  DeDe achieves parallelism by decomposing the full problem into per-resource and per-demand allocations. Therefore, when we cannot fully decompose the problem, DeDe will experience limited parallelism, although this does not happen in our survey (Table 1).
@@
Suboptimal solutions.  It is possible for DeDe not to attain an ideal solution when the allocation problem deviates from the standard ADMM formulation.
  • Non-convex problems.  ADMM can solve non-convex optimization problems [53, 62, 68] empirically, providing theoretical guarantees only under specific assumptions[56, 38, 60]. As a result, DeDe can output suboptimal solutions in certain non-convex cases.
@@ 5 Case Studies
In this section, we will present case studies on how DeDe can be applied to three resource allocation problems: traffic engineering, cluster scheduling, and load balancing, as also studied in POP [43].
@@ 5.1 Traffic Engineering
  • Objective.  The utilization of a link $(u,v)$ is the ratio of the total flow assigned to the link to its capacity. Our objective is to minimize the max link utilization across the network:
@@
Using DeDe.  The traffic engineering problem aligns the separable structure for DeDe: the objectives are the sum of allocation quality per demand or the maximum of allocation quality per link, respectively. Linear constraints are applied to each demand and each link. Therefore, DeDe can decompose the problem into $|E|$ per-link allocations and $|V|^{2}$ per-demand allocations for parallelism.
@@ 5.2 Cluster Scheduling
Using DeDe.  The cluster scheduling exhibits the separable structure required by DeDe: The objective function is a minimum of allocation quality per job and a sum of allocation quality per job, respectively. Additionally, linear constraints are applied to each job and each computing resource. Thus, DeDe can decompose the problem into $n$ per-job allocations and $m$ per-demand allocations.
@@ 5.3 Load Balancing
Using DeDe.  Given the separable structure, where the objective is the sum of allocation quality per data shard and linear constraints apply to each server, and each shard, DeDe can decompose the problem into $n$ per-server allocations and $m$ per-shard allocations.
6 Implementation of DeDe
We have implemented DeDe as a Python package and uploaded it to PyPI for easy installation: pip install dede.
Listing 1 showcases a simple resource allocation problem solved using our DeDe package. In line 4, we initially create a $N\times M$ non-negative allocation variable x, where x[i,j] represents the allocation of resource i to demand j. We then create a constraint for each resource in line 7–8 and a constraint for each demand in line 9–10, as well as an objective of minimizing the sum of x in line 13. Finally, we construct the problem in line 16–17 and solve it in line 18.
1import dede
@@ 6 Implementation of DeDe
3# Create allocation variables
4x = dede.Variable((N, M), nonneg=True)
@@
12# Create an objective
13obj = dede.Minimize(x.sum())
@@
15# Construct and solve the problem
16prob = dede.Problem(
@@
Listing 1: Resource allocation example with DeDe package.
DeDe borrows the interface from cvxpy [17, 4]. DeDe inherits most of cvxpy’s methods, e.g. Variable(.) in line 4, Minimize(.) in line 13. One difference from cvxpy is that users must specify resource constraints and demands constraints separately (line 16–17) when initializing a problem. As a parallel solution, DeDe allows users to specify the number of CPU cores (line 18); if not specified, DeDe uses all available CPU cores.
Behind DeDe’s interface, the resource allocation problem is solved in the following three steps.
@@ 7 Evaluation
    How does DeDe compare with state-of-the-art approaches in allocation quality and computation time? (§7.1)
@@
    How do DeDe and state-of-the-art approaches react to changes in problem granularity, temporal fluctuation, spatial redistribution, and link failures? (§7.2)
@@
    What is the contribution of each design in DeDe? (§7.3)
@@
  • DeDe.  DeDe is our truly parallel implementation, with details described in Section 6.
  • DeDe.  To facilitate a fair comparison with POP, we create a variant of DeDe, denoted as DeDe, following POP-style simulation where subproblems are solved sequentially and a parallel solving time is calculated mathematically.
@@
7.1 DeDe vs. the State of the Art
In this section, we compare DeDe against state-of-the-art approaches on resource allocation problems from Section 5.
@@ 7.1.1 Traffic Engineering
Maximize total flow.  Figure 3 compares DeDe to state-of-the-art methods in maximizing total flow. DeDe achieves 90.8% satisfied demand within 30 seconds and 92% within 60 seconds, demonstrating a superior trade-off between satisfied demand and computation time compared to Exact Sol, POP, and Pinning. While POP-4 also reaches 92% satisfied demand, its average computation time is significantly longer at 1658 seconds. POP-16 and POP-64 improve parallelism by dividing the workload into more subproblems, reducing computation times to 456 seconds and 380 seconds, respectively. Note that the speedup of POP-64 is limited compared to POP-16 as the number of cores per subproblem drops from 4 to 1. This increased parallelism also comes at a cost, with satisfied demand decreasing to 87.4% for POP-16 and 81.6% for POP-64. Pinning focuses on top demands to reduce problem size but remains constrained by the sequential nature of its optimization solvers. As a result, it achieves a modest speedup, requiring 149 seconds to reach 87.3% satisfied demand. Teal, a learning-accelerated approach, leverages massive GPU parallelism to achieve 89% satisfied demand in just 1 second. However, it relies on a highly customized machine-learning framework, demanding significant human effort for design and training.
DeDe, representing the same algorithm under ideal conditions as POP, achieves a satisfied demand that is 8% higher than DeDe within 10 seconds in Figure 3 and consistently delivers superior allocation quality across the other figures. The reason is that DeDe executes each iteration more quickly, allowing it to complete more iterations within a given time, resulting in higher allocation quality. The difference in iteration time between DeDe and DeDe arises from three factors, as discussed in more detail in § 7.3: (1) DeDe measures the end-to-end time, including problem-solving, compilation, solution unpacking, and other associated tasks, while DeDe only takes the problem-solving time into account. (2) DeDe’s parallel implementation (§ 6) incurs overhead from cache contention, which slows down the problem-solving process. (3) In DeDe, each subproblem is statically pre-assigned to one of the processes, making it susceptible to stragglers. In contrast, DeDe assumes perfect dynamic scheduling.
@@
Min-max link utilization.  Figure 4 compares DeDe with other approaches in minimizing maximum link utilization. DeDe achieves a low maximum link utilization of 1.67 within 10 seconds and 1.71 within 5 seconds. Exact Sol reaches 1.63 with an average computation time of 35 seconds, and it is faster than itself in maximizing total flow because minimizing max link utilization is "easier" and requires fewer iterations. Pinning takes 42.5 seconds, making it slower than DeDe and slightly slower than Exact Sol, as it requires additional time to rebuild the problem when the top 10% of node pairs change. POP’s variants, POP-4, POP-16, and POP-64, achieve a maximum link utilization of 1.70, 1.77, and 1.95 with 9, 2, and 0.4 seconds on average, respectively. Although POP-16 is faster than DeDe, a fairer comparison between POP-16 and DeDe, both simulating parallelism, reveals that DeDe achieves a comparable computation time to POP-16. The domain-specific solution Teal takes only 0.3 seconds with massive parallelism from GPU with a maximum link utilization of 1.74.